OpenStudy (anonymous):

Solve for "x" 2(3^x)=12^x-1.

OpenStudy (ahsome):

Is this it? \[2\times3^x=12^{x-1}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

this one takes a minute gimmick is to write \[12=3\times 4\] so \[12^{x-1}=3^{x-1}4^{x-1}\] then divide both sides of \[2\times 3^x=3^{x-1}4^{x-1}\] by \(3^{x-1}\)

OpenStudy (anonymous):

maybe @Ahsome has a slicker way, not sure if that method is best

OpenStudy (anonymous):

maybe adding a log will help?

OpenStudy (ahsome):

Hmmm..

OpenStudy (ahsome):

Would we use logarithim?

OpenStudy (anonymous):

i believe so

OpenStudy (anonymous):

once you ave \[6=4^{x-1}\] you would use the change of base formula so yes, logarithms for sure

OpenStudy (anonymous):

but that is the last step, not the first step

OpenStudy (anonymous):

how did you get there?

OpenStudy (ahsome):

According to wolframalpha, \[x\approx2.2925\]

OpenStudy (anonymous):

i wrote the steps above take a look and see if you have any questions about how i got to \[6=4^{x-1}\]

OpenStudy (ahsome):

\[2×3^x=12^{x−1}\] \[3^x=\frac{12^{x−1}}{2}\] \[log_3{\frac{12^{x−1}}{2}}=x\]

OpenStudy (ahsome):

Hmmm

OpenStudy (anonymous):

which you and also write as \[24=4^x\] making \[x=\frac{\ln(24)}{\ln(4)}\]

OpenStudy (anonymous):

yeah i get that

OpenStudy (ahsome):

Wait a minute is: \[2×3^x=12^{x−1}\] The same as: \[2×3^x=12^{x}\div12^1?\]

OpenStudy (anonymous):

How did you go from 12=3×4 to 12^x-1=3^x-14^x-1

OpenStudy (ahsome):

Me, or @satellite73?

OpenStudy (ahsome):

\[2×3^x=12^{x−1}\] \[2×3^x=12^x÷12^1\] \[12(2\times3^x)=12^x\] \[log_{12}{12(2\times3^x)}=x\]